OpenStudy (anonymous):

2x^2+4x=8 For x please?

OpenStudy (lalaly):

2x^2+4x-8 = 2(x^2+2x-4) use the quadratic formula

jimthompson5910 (jim_thompson5910):

Solutions are \[\large x=-1+\sqrt{5}\] or \[\large x=-1-\sqrt{5}\] See attached to see how I got those answers.